### v1

# Infra-red thermographic inversion in ST40

Abstract
A new tool for infra-red thermographic inversion on Tokamak Energy’s spherical tokamak, ST40, is developed in-house and here presented. Functional Analysis of Heat Flux (FAHF) is written in Python, and configured for multi-2D thermographic inversions, solving the heat conduction equation within the divertor tiles via the finite difference method, and leveraging an explicit time-marching scheme. Using infra-red camera data of the highest effective spatial resolution available, FAHF ultimately outputs the plasma perpendicular heat flux on the divertor, crucial quantity in any edge plasma investigation. In the present work, the internal numerics of the tool is first successfully verified by formal time and space convergence analyses, and corroborated by an energy balance assessment. A significant sensitivity of the perpendicular heat flux computed by FAHF to the user-selected spatial resolution is then evidenced. However, a precise heat flux is proved to be recoverable by ensuring a sufficiently high resolution. Last, the appropriateness of the model/geometry simplifications adopted in FAHF is successfully confirmed, by means of comparison against COMSOL Multiphysics® simulations. FAHF is hence conclusively proven to qualify as a precise and accurate tool for infra-red thermographic inversions.
Keywords: ST40, infra-red thermography, thermographic inversion, heat flux, scrape-off layer plasma, finite difference method.
Introduction
Tokamak-based nuclear fusion presents a plethora of physico-technological challenges to be overcome for its achievement [1]. A robust power exhaust (PEX) solution figures among them. This particular challenge arises from the substantial heat the fusion-grade core plasma leaks in the peripheral region of open magnetic field lines, the scrape-off layer (SOL). This power, expected to amount to hundreds of MW in a thermonuclear reactor [2, 3], is unusable by the plasma to generate further fusion reactions, and it is invariably channelled by the SOL out of the plasma system, towards the plasma-facing components (PFCs).
However, irrespective of the total available area offered by the PFCs, the peculiar anisotropicity of the SOL plasma transport [4, 5] favours a severely localised heat deposition [6], which nominally translates in heat flux densities in the order of few hundreds of $\text{MW}\cdot\text{m}^{-2}$ to be handled. Although mitigatory action is routinely undertaken in tokamaks [7, 8, 9], the baseline PFC technology currently available and tested has successfully demonstrated[10] suitable PEX properties and mechanical integrity only up to few tens of $\text{MW}\cdot\text{m}^{-2}$. Carefully determining the value attained by the parameters affecting the heat flux footprint hence becomes vital. Among them are the peak heat flux $q_{0}$ and power fall-off length $\lambda_{q}$, respectively dictating the severity of the flux, and the width of the footprint [6].
The well-established diagnostic method of infra-red (IR) thermography [11] offers the possibility of a detailed determination of such parameters, and of a thorough multi-dimensional characterisation of the SOL plasma at the interface with the PFCs. By providing a two-dimensional (2D) view of the thermally loaded PFCs, the evolution in time of the PFC surface temperature diagnosed by the IR camera is leveraged to infer the heat flux which caused such evolution to occur. This inference is known as thermographic inversion, and involves the solution of the heat conduction (Fourier) equation [12]. But, among others, the three-dimensionality of the PFCs [13], surface imperfections and deposited impurity layers[14, 15] significantly exacerbate the modelling complexity. A non-exhaustive list of teams reporting on IR thermographic apparati and/or thermographic inversion analyses which aids further insights is: Alcator C-Mod[16], ASDEX[17, 18, 19], COMPASS[20], DIII-D[21], EAST[22], ITER[23], JET[24, 25], KSTAR[26, 27], LHD[28], MAST[29, 30], NSTX[31], SPARC[32, 33], TCV[34, 35], W7-X[36], and WEST[37, 38, 39].
In the current work, the new in-house Python-based tool for the Functional Analysis of Heat Flux (FAHF) in the ST40 spherical tokamak [40, 41], owned and operated by Tokamak Energy Ltd., is presented. With the goal of constituting a fast and flexible inter-shot analysis tool, simplifications are accordingly adopted in FAHF. Therefore, in addition to the customary code verification to quantify the code’s precision, and assessment of the output physical parameters, the appropriateness of the simplifying assumptions of FAHF is scrutinised by means of comparison with the renowned COMSOL Multiphysics® modelling suite [42], to estimate its accuracy. Cross-code comparisons indeed constitute an effective method for the quantification of model and geometry errors [43, 44], paramount in the present context.
The paper is organised as follows. Sec. 2 expounds on the ST40 tokamak, experimental setup and data selection. The numerical methods and tools adopted are described in Sec. 3. The results of the study are collated in Sec. 4, and discussed in Sec. 5. Lastly, Sec. 6 summarises the conclusion of the work, and comments on future studies.
Experimental setup
The ST40 tokamak
ST40 is a high-field compact spherical tokamak (ST) constructed and operated by Tokamak Energy Ltd., characterised by a major radius $R=0.4-0.5$ m, minor radius $a=0.2-0.25$ m, and a magnetic field up to 2.0 T on axis.
Disconnected double-null diverted plasmas are most commonly featured by ST40. Fig. 1a shows the EFIT magnetic reconstruction [45] in the poloidal cross section for shot 11376 at time instant 128 ms (Sec. 2.3).
ST40 is provided with highly-shaped fish-scaled[13] tiled divertor targets, pictured in Fig. 1b, with inter-tile gaps of 4.5 mm in width. The leading edge of each tile is poloidally chamfered[13], an area that is not of interest in the present study, and therefore neglected. The leftover volume of the divertor tiles is thus a rectangular cuboid, with 4 mm of plasma-facing TZM molybdenum (Mo) overlaid to 25 mm of copper-cromium-zirconium (CuCrZr). Poloidally, the geometry of the divertor is that of an open non-baffled divertor, with the inner leg orthogonal to the plate, and an horizontal outer leg (Fig. 1a).
The reader is referred to Ref. [41] for further information about ST40, and to Ref. [46] for specifics of its divertor.
Diagnostic apparatus: IR camera
ST40 is equipped with a FLIR X6903sc MWIR ($3-5\;\mu$m) camera, which features an acquisition rate of 0.8 kHz ($\Delta t_{\text{IR}}=$ 1.25 ms) for its 612 $\times$ 540 pixels, and diagnoses the upper outer divertor target (Fig. 1a). The camera was installed in collaboration with Oak Ridge National Laboratory. Its line of sight (LoS) on the divertor provided by an ad-hoc re-entrant optical endoscope is close to perpendicular to the surface, and the field of view (FoV) covers the minimum amount of surface required to fully capture the heat flux footprint. The compenetration of the above LoS and FoV hence maximises the camera effective spatial resolution along the divertor, which results in a noteworthy $\Delta x_{\text{IR}}=0.21\pm 0.03$ mm/pixel ($\sim 5$ pixel/mm), appreciable in Fig. 1c. The comparison with the list of references in Sec. 1 actually suggests the spatial resolution of ST40 IR camera to be the highest available worldwide for thermographic inversions of a full plasma heat flux footprint.
The calibration of the IR camera and endoscope has been accomplished ex-situ via a known black-body emission source, in the temperature window $[20;118]$. The surface emissivity of the divertor tiles is instead assumed to be unitary, constant with respect to the position in space and surface temperature. This does not impact the main focus of the present work.
Figure 1: (a) EFIT reconstruction of the disconnected double-null magnetic equilibrium of ST40’s shot 11376 at 128 ms: primary separatrix in magenta solid; secondary separatrix in magenta dashed; further magnetic surfaces in grey; wall in black; chord 10 in blue (1:1 scale). (b) IR camera view at 128 ms on the upper outer divertor in pixel space $(x_{\text{p}},\phi_{\text{p}})$. Overlaid are (see text for details): the FAHF reference frame in real space with $\phi$, proxy for toroidal direction, $x$, proxy for poloidal direction, and the direction in the depth of the tiles $y$ (entering the screen); an illustrative representation of chord 10 of the bundle in both its sub-chords (one per tile, blue dots). (c) Temperature $T_{\text{IR}}$ as extracted from the IR view via the Calcam software [47] (Appendix A) along chord 10 in real space, with the empty gap in its middle. The data-point resolution is the actual $0.21\pm 0.03$ mm/pixel of the ST40 IR camera.
Reference shot
Shot number 11376 within the ST40 database is selected for the analyses of the present study. Reasons behind the choice are primarily the availability and quality of the IR experimental data.
Furthermore, shot 11376 is a representative instance of the ”average” ST40 plasma shot, from the edge plasma point of view. It is in fact a quiet L-mode, featuring a single-exponential heat flux footprint (Sec. 4.1.3).
Additionally, during the time window $[50;140]$ ms the plasma presents a stable flat-top. Although the IR camera is marginally affected by saturation in $[130;140]$ ms in the region of interest (Sec. 3.3.1), this does not impact the numerical characterisation of FAHF.
Despite the magnetic equilibrium of shot 11376 nominally being a lower-biased disconnected double-null (Fig. 1a), the low degree of disconnection ($\delta r_{\text{sep}}/\lambda_{q}\sim 1.5$ at the outer mid-plane [48]) allows the plasma to deposit a satisfactory amount of power in the camera FoV during the actual divertor heating phase in $[90;140]\;\text{ms}$.
Numerical methods and tools
FAHF
Theoretical fundamentals and model assumptions
Safely discarding any volumetric heat source/sink (e.g. arising from induced currents), the general form of the unsteady homogeneous heat conduction equation for an isotropic medium in cartesian coordinates reads [12]:
$\varrho c_{\text{p}}\frac{\partial T}{\partial t}=\nabla\cdot(\kappa\nabla T),$ (1)
with $\varrho$ being the mass density ($\text{kg}\cdot\text{m}^{-3}$), $c_{\text{p}}$ the specific heat capacity at constant pressure ($\text{J}\cdot\text{kg}^{-1}\cdot\text{K}^{-1}$), and $\kappa$ the heat conductivity ($\text{W}\cdot\text{m}^{-1}\cdot\text{K}^{-1}$).
Crucially, if $\kappa=\kappa(T)$ or it explicitly depends on space (e.g. multi-material PFC), then $\nabla\cdot(\kappa\nabla T)=\nabla\kappa\cdot\nabla T+\kappa\nabla^{2}T$, and a careful differentiation would be required. An elegant solution to this is employed in Ref. [49], which however forces an analytical form for the temperature-dependent thermal properties to be assumed. But for the specific case of ST40, the divertor heating phase during the plasma flat top lasts around 50 ms at most ($[90;140]\;\text{ms}$, Sec. 2.3). Therefore:
  • Assumption #1: the material properties are assumed to be temperature-independent.
  • Assumption #2: the divertor tiles are assumed to be a single layer of Mo of 29 mm in thickness, instead of the actual 4 mm of plasma-facing Mo overlaid to 25 mm of CuCrZr.
With respect to Asm. #2, the region affected by a temperature variation in a semi-infinite solid [^1] With $L=29$ mm of the present case, a reasonably constant heat flux over $[90;140]$ ms (Fig. 6), and assuming $\text{Fo}=\alpha t/L^{2}<0.2$ [12], the approximation of semi-infinite solid would hold for up to $\sim 3$ s. is here indeed limited to $2\sqrt{\alpha\cdot 50\;\text{ms}}\sim 3.3$ mm (with $\alpha\sim 54\;\text{mm}^{2}\cdot\text{s}^{-1}$ defined in Eq. (2)) along the depth of the tile, hence still confined within the only Mo stratum. No impurity surface layer [14, 49] is accounted for.
Asms. #1 and #2 imply model errors, which are formally assessed in Sec. 4.2.2. Eq. (1) accordingly reduces to:
$\frac{\partial T}{\partial t}=\alpha\nabla^{2}T,$ (2)
where $\alpha=\kappa/(\varrho c_{\text{p}})$ is the thermal diffusivity ($\text{m}^{2}\cdot\text{s}^{-1}$), unaffected by any differential operator only because of its independence on $T$ in the present case.
Dimensionality of the problem and geometrical assumption
The orthogonal reference frame adopted in FAHF is pictured in Fig. 1b and chosen such as to be aligned with the sides of the tiles. In particular, $x$ is the direction along the divertor (henceforth referred to as ”poloidal direction”), $y$ the direction along the depth of the tile, and $\phi$ is orthogonal to $x$ and $y$, and happens to be a proxy for the toroidal direction (which would actually be aligned to the strike point band). In ST40:
  1. 1.
    Assumption #3: the two-dimensional description of heat conduction is assumed to be appropriate, and the toroidal direction is neglected by assuming toroidal symmetry.
A bundle of $N_{\phi}$ independent parallel observation chords is generated, one such being shown in Fig. 1b (blue dots). The chords are aligned with the $x$ direction and sit at uniformly-spaced locations $\{\phi_{k}\}_{k=1}^{N_{\phi}}$ on the rectangular planar surface (Sec. 2.1) of the divertor tile CAD ($x\phi$ plane). This CAD is a suitable representation of the actual tiles, as long as surface imperfections are absent. Surface intrusions/protrusions deterministically cause noticeable abnormalities in the results. Any chords affected is therefore discarded from the analysis.
Each chord is then further divided along $x$ in two equal sub-chords, one per tile, where the gap running along $\phi$ lies. This is allowed by the thermal insulation provided by the gap (Sec. 3.1.4). The resulting 2D computational domain in FAHF is therefore a (32 mm) $\times$ (29 mm) rectangle in the $xy$ plane. The overall output along the entire chord (e.g. surface heat flux profile, Sec. 3.1.6) is ultimately obtained by juxtaposing in space the FAHF output of each of the two sub-chords at every time-step, as accomplished for $T_{\text{IR}}$ in Fig. 1c.
Asm. #3, relying on FAHF being independently run along (each sub-chord of) each chord, is justified by the fact that the poloidal temperature gradients $\partial_{x}T_{\text{IR}}$ dominate over the toroidal gradients $\partial_{\phi}T_{\text{IR}}$ when the strike point lies in the camera FoV, i.e. the time window of most interest $[90;140]\;\text{ms}$. However, following from the peculiar three-dimensionality of the ST40 divertor, $\partial_{\phi}T_{\text{IR}}\neq 0$ holds, as evident from Fig. 1b. Therefore, although FAHF’s multi-2D method is still preferred over a full 3D treatment for reasons of computational cost, the geometrical error implied by the multi-2D approach is assessed in Sec. 4.2.1.
Numerical discretisation in space and time
Following from the dimensionality of which in Sec. 3.1.2, and the unsteady Eq. (2), FAHF is configured as a 2D time-dependent heat conduction solver in cartesian coordinates ($xy$ plane). The finite difference (FD) method is relied upon for the spatial discretisation in the nodes of the rectangular domain, for reasons of simplicity and speed.
In particular, $N_{x}$ points $\{x_{i}\}_{i=1}^{N_{{x}}}$ are placed along each sub-chord ($x$ direction), with uniform spacing $\Delta x=x_{i+1}-x_{i}\;\forall i\in[1;N_{x})$. Instead, $N_{y}$ points $\{y_{j}\}_{j=1}^{N_{{y}}}$ are distributed along the tile depth ($y$ direction), with uniform spacing $\Delta y=y_{j+1}-y_{j}\;\forall j\in[1;N_{y})$ (potentially differing from $\Delta x$), where $j=1$ identifies the plasma-facing surface, and $j=N_{y}$ the back-surface. The second-order centred FD scheme[12] is then adopted in the interior nodes $(i,j)\in[2;N_{x}-1]\times[2;N_{y}-1]$.
In time, the explicit first-order forward Euler time-marching scheme is implemented at the time instants $\{t^{n}\}_{n=1}^{N_{t}}\in[0;200]$ ms, with uniform time-step $\Delta t=t^{n}-t^{n-1}\;\forall n\in(1;N_{t}]$. This implementation must satisfy the stability criterion [12]:
$\Delta t\cdot\left(\frac{1}{\Delta x^{2}}+\frac{1}{\Delta y^{2}}\right)<\frac{1}{2\alpha}.$ (3)
The temperature at the node $(i,j)$ and time instant $n$ is finally denoted as $T_{i,j}^{n}=T(x_{i},y_{j},t^{n})$.
Boundary and initial conditions
At the plasma-facing boundary of the computational domain ($j=1$), a Dirichlet-type boundary condition (BC) is enforced via the experimental temperature data. Exploiting the Calcam software[47] according to Appendix A, $T_{\text{IR}}(x,\phi)$ is sampled over the set $\{x_{i}\}_{i=1}^{N_{x}}$ along each chord at every time instant $\{t^{n}\}_{n=1}^{N_{\text{t}}}$, obtaining the BC data $T_{i,1}^{n}=T_{\text{IR},i}^{n}\;\forall i,n$. Although both the user-selected $\Delta x$ and the $\Delta t$ resulting form the stability criterion (Eq. (3)) might fall below the IR camera spatial and/or temporal resolution (Sec. 2.2), this is not here of concern, as the linear interpolation of $T_{\text{IR}}$ is checked to not introduce any artifacts (e.g. over-/under-shooting).
On the remaining non-plasma-facing boundaries ($\{i=1\}\cup\{i=N_{x}\}\;\forall j>1$ or $j=N_{y}\;\forall i$), first-order homogeneous Neumann BCs are enforced. Physically, homogeneous Neumann BCs represent adiabatic (no-flux) boundaries, and implicitly assume the sub-chords (divertor tiles) to be thermally independent ($\{i=1\}\cup\{i=N_{x}\}\;\forall j>1$). This holds under the legitimate assumptions of a negligible black-body emission in the gap ($T_{\text{IR}}<50$), and of a negligible heat convection/conduction provided by the rarefied neutral gas filling this volume (few Pa in pressure at most [50]). Additionally, the thermal contact of the tile back-surface ($j=N_{y}\;\forall i$) with the supporting structures is also neglected, according to Sec. 3.1.1 -- the variation of the bulk temperature anywhere more than few millimetres away from the plasma-facing boundary being negligible.
In the temporal domain, the initial condition (IC, $t=0$ ms) $T_{i,j}^{1}$ chosen at the first time instant $n=1$ is a uniform distribution in the entire domain. Its value is the spatial-average temperature along the sub-chord as measured by the IR camera, i.e. $T_{i,j}^{1}=\sum_{i=1}^{N_{x}}T_{\text{IR},i}^{1}/N_{x}\;\forall i,j$. Different flavours of such IC are tested, not providing any appreciable difference in the outcome after a few time-steps.
Discretised form of the problem
The resulting discretised form of the heat conduction problem FAHF solves throughout the entire plasma discharge ultimately reads:
$\begin{split}&\begin{split}\frac{T_{i,j}^{n}-T_{i,j}^{n-1}}{\Delta t}&=\alpha\cdot\frac{T_{i+1,j}^{n-1}-2T_{i,j}^{n-1}+T_{i-1,j}^{n-1}}{\Delta x^{2}}+\\ &+\alpha\cdot\frac{T_{i,j+1}^{n-1}-2T_{i,j}^{n-1}+T_{i,j-1}^{n-1}}{\Delta y^{2}}+\\ &+o(\Delta x^{2})+o(\Delta y^{2})+o(\Delta t).\end{split}\\ \\ &\begin{split}&\text{IC: }T^{1}_{i,j}=\sum_{i=1}^{N_{x}}T_{\text{IR},i}^{1}/N_{x}\text{ if $n=1\;\forall i,j$.}\\ &\text{BC: }\text{Dirichlet }T^{n}_{i,1}=T_{\text{IR},i}^{n}\text{ if $j=1\;\forall i,n\geq 2$.}\\ &\begin{split}\text{BC: }&\text{homogeneous Neumann (no-flux)}\\ &\text{if $\{i=1\}\cup\{i=N_{x}\}\;\forall j>1,n\geq 2$}\\ &\text{or $j=N_{y}\;\forall i,n\geq 2$.}\end{split}\end{split}\end{split}$ (4)
The only unknown $T_{i,j}^{n}\;\forall n\in[2;N_{t}]$ is explicitly retrieved at each timestep, and embodies the output of the simulation with a discretisation error $o(\Delta x^{2})+o(\Delta y^{2})+o(\Delta t)$.
Surface heat flux and Eich fit
The quantity of primary physical interest of the IR thermographic inversion is the surface heat flux $q\;(\text{W}\cdot\text{m}^{-2})$ perpendicular to the PFC surface (i.e. along $y$). From the temperature distribution $T_{i,j}^{n}$ computed by the algorithm via Eq. (4), the surface heat flux is computed in post-processing at the plasma-facing surface ($j=1$) according to the first-order discretisation:
$\begin{split}q_{i}^{n}=&-\kappa\cdot\frac{T_{i,2}^{n}-T_{i,1}^{n}}{\Delta y}+\\ &+o(\Delta x^{2})+o(\Delta y)+o(\Delta t),\end{split}$ (5)
with $q_{i}^{n}=q_{i,1}^{n}$. The error $o(\Delta x^{2})+o(\Delta t)$ is inherited from $T_{i,j}^{n}$ of Eq. (4), while $o(\Delta y^{2})\ll o(\Delta y)$ by definition. Higher-order spatial discretisations[49] are available, but here not required.
Once $q_{i}^{n}$ is known for two adjacent sub-chords, fitting of the juxtaposed profile (Sec. 3.1.2) is accomplished by leveraging the Eich function $q_{\text{Eich}}$ from Eq. (1) of Ref. [6]. Although an iterative procedure is actually best-posed in presence of the multi-parameter non-linear fitting function[46] $q_{\text{Eich}}$, this is not here accomplished for the sake of simplicity. Instead, the fit is deemed satisfactory upon showing $R^{2}>98.5\%$. This simplified approach still allows to convey an equivalent message.
Ultimately, the main output of the fitting procedure is embodied by the physical parameters of: maximum heat flux $q_{0}$, power fall-off length $\lambda_{q}$, spreading factor $S$, background heat flux $q_{\text{BG}}$, strike point location $x_{0}$, and poloidal flux expansion $f_{X}$. The reader is redirected to Ref. [6] for detailed descriptions of such quantities. The integral value of the Eich profile $q_{\text{Eich}}(x)$ is here additionally added, and defined as:
[…]
Internal energy variation and incoming energy
Sections 3.1.5 and 3.1.6 provide the building blocks to compute the internal energy variation (per unit length, along the non-modelled $\phi$ direction) throughout the volume of the domain:
$\begin{split}\Delta\mathcal{E}_{\text{int}}^{n}&=\sum_{i=1}^{N_{x}}\sum_{j=1}^{N_{y}}\varrho c_{\text{p}}(T_{i,j}^{n}-T_{i,j}^{n-1})\cdot\Delta x\Delta y+\\ &+o(\Delta x^{2})+o(\Delta y^{2})+o(\Delta t),\end{split}$ (7)
and the net energy (per unit length) flowing through the plasma facing boundary:
$\begin{split}\mathcal{E}_{\text{surf}}^{n}&=\sum_{i=1}^{N_{x}}q_{i}^{n}\cdot\Delta x\Delta t+\\ &+o(\Delta x^{2})+o(\Delta y)+o(\Delta t),\end{split}$ (8)
which enable the characterisation of the energy-preserving properties of FAHF. In Eqs. (7) and (8) the errors follow from Eqs. (4) and (5), respectively.
COMSOL Multiphysics®
COMSOL Multiphysics® [42], herein abbreviated to ”COMSOL”, is a widely-adopted state-of-the-art tool for multiphysical modelling. This suite is here configured as a pure heat conduction solver, and therefore utilises the finite element method (FEM) in space, with quadratic Lagrange elements, and second-order backward time stepping [42].
A variety of different models, detailed in Sections 4.2.1 and 4.2.2, is considered in COMSOL. This is an attempt of matching FAHF the closest in dedicated instances, and quantifying the model/geometry errors of FAHF in others, ultimately resulting in an estimate of its accuracy.
The same computational domain of FAHF (Sec. 3.1.2) is uniformly meshed with $N_{x}\times N_{y}=100\times 100$ rectangular elements ($\Delta x=3.5\cdot 10^{-4}$ m and $\Delta y=2.9\cdot 10^{-4}$ m), and the same initial and boundary conditions (Sec. 3.1.4) are adopted, with $\Delta t=1\cdot 10^{-4}$ s. Because of reasons related to ingent computational cost, only one divertor tile (bottom left in Fig. 1b) is modelled by COMSOL in 3D and, for consistency, only the corresponding sub-chord of chord 10 (Sec. 3.3.1) belonging to such tile is modelled in 2D. Clearly, this time-saving choice does not impact any of the conclusions inferred.
Ancillary convergence tests of the COMSOL results have been performed, but are not here reported. Noteworthy is instead the need of de-activating the automatic smoothing of the resulting heat flux in COMSOL, which would otherwise constitute a further unwanted source of error, for not being implemented in FAHF.
Definitions and conventions
Reference chord
Chord number 10 ($k=10$, Sec. 3.1.2) is taken as a reference, and depicted in Fig. 1b (blue dots). This chord sits sufficiently far away from poloidal-running gaps to not show any deleterious boundary effect [46]. Also, chord 10 is affected neither by magnetic shadowing [46] nor by surface imperfections (Sec. 3.1.2), both which would spuriously distort the heat flux pattern.
Throughout the following, only chord 10 is investigated, but similar results apply for all the other ones. The output of a representative multi-chord analysis is reported by Ref. [46].
Convergence error
The behaviour of the errors in time $\varepsilon_{*}(\Delta t)$ and space $\varepsilon_{*}(\Delta x)$ is assessed via time and space convergence, the latter forcing $\Delta x=\Delta y$ for the sake of simplicity. These procedures are a numerical characterisation of the precision of the algorithm (verification), ensure trust in the internal workings of the code, and suggest the most suitable choice of $\Delta t$ and $\Delta x$, within the limit posed by Eq. (3). These processes only involve the actual output of the simulation, i.e. the temperature distribution $T_{i,j}^{n}$ (Sec. 3.1.5), and are agnostic of the heat flux $q_{i}^{n}$ (Sec. 3.1.6).
In particular, in the time convergence, the time-step $\Delta t$ is varied in the interval $[4\cdot 10^{-6};4\cdot 10^{-4}]$ s. The actual values are detailed in Table 1. Instead, Table 2 reports the values $\Delta x\in[6.8\cdot 10^{-5};2.3\cdot 10^{-3}]$ m of the space convergence. Linear interpolation is employed whenever $\Delta t<\Delta t_{\text{IR}}=1.25\cdot 10^{-3}$ s or $\Delta x<\Delta x_{\text{IR}}=2.1\cdot 10^{-4}$ m (Sec. 2.2).
Each simulation in the time convergence is run with a fixed spatial discretisation with $\Delta x=5.5\cdot 10^{-4}$ m. This space-step is small enough such that to not alter the error in time across all the cases, i.e. $o(\Delta x^{2})\ll o(\Delta t)\;\forall\Delta t$ in the time convergence. In the space convergence, $\Delta t=7.2\cdot 10^{-6}$ s dually guarantees $o(\Delta t)\ll o(\Delta x^{2})\;\forall\Delta x$.
Common to both procedures, a reference point in time $t_{*}=128$ ms and a reference point in space $(x_{*},y_{*})=(16.55;2.21)$ mm are selected where to locally compare the temperature across the various discretisations.
The time point $t_{*}=128$ ms lies well within $[90;140]\;\text{ms}$, the time window of most interest (Sec. 2.3). Unless otherwise specified, the results obtained at $t=t_{*}$ apply to any other time instant. The point $(x_{*},y_{*})$ lies in the volume of the domain, within $\sim 3$ mm of the plasma-facing surface so as to be appreciably affected by a temperature variation (Sec. 3.1.1), and away from any other boundaries, making it suitable for comparison across different cases. This ”local” approach is preferred over integral errors (in the fashion of Sec. 3.3.3) to minimise any near-boundary artifacts which would impair a precise convergence analysis.
The reference temperature $\tilde{T}_{*}=\tilde{T}(x_{*},y_{*},t_{*})$ to compare the other cases against is defined as the result obtained with the smallest time/space-step of the time/space convergence, i.e. the most precise instance. The relative difference of the resulting $T_{*}=T(x_{*},y_{*},t_{*})$ of a given simulation with respect to the reference is labelled as ”error” and it is computed according to:
$\varepsilon_{*}=\frac{\left|T_{*}-\tilde{T}_{*}\right|}{\left|\tilde{T}_{*}\right|}.$ (9)
Clearly, the error of the very reference case is zero by definition, and it is therefore discarded.
Heat flux error
With regards to the surface heat flux (Sec. 3.1.6), the relative difference of the result $q_{i}^{n}$ of a given simulation (e.g. COMSOL, Sec. 3.2) with respect to FAHF’s $\tilde{q}_{i}^{n}$ is referred to as ”heat flux error” $\varepsilon_{q}$. Such error leverages the $\ell^{1}$ norm, with the heat fluxes being interpolated in time and space onto the grid of the most refined case between the two.
The error can be computed either via integration in space:
$\varepsilon^{n}(q)=\frac{\sum_{i=1}^{N_{x}}\left|q_{i}^{n}-\tilde{q}_{i}^{n}\right|}{\sum_{x=1}^{N_{x}}\left|\tilde{q}_{i}^{n}\right|},$ (10)
returning an error profile in time or, viceversa, via integration in time:
$\varepsilon_{i}(q)=\frac{\sum_{n=1}^{N_{t}}\left|q_{i}^{n}-\tilde{q}_{i}^{n}\right|}{\sum_{n=1}^{N_{t}}\left|\tilde{q}_{i}^{n}\right|},$ (11)
returning an error profile in space along $x$ (at the plasma-facing surface). This ”integral” approach allows accounting for all the areas of agreement/disagreement among two cases.
Energy balance error
Starting from Sec. 3.1.7, the assessment of the energy conservation in each case of time and space convergence is accomplished. Similarly to Sec. 3.3.2, only the time instant $t_{*}=128$ ms is considered in the error estimation.
For a perfect energy balance to be satisfied, the internal energy variation (Eq. (7)) must match the net incoming energy (Eq. (8)). The relative difference among $\Delta\mathcal{E}_{\text{int*}}=\Delta\mathcal{E}_{\text{int}}(t_{*})$ and $\mathcal{E}_{\text{surf*}}=\mathcal{E}_{\text{surf}}(t_{*})$ for each case, labelled as ”energy balance error”, is thus defined by:
$\varepsilon_{*}(\mathcal{E})=\frac{\left|\Delta\mathcal{E}_{\text{int*}}-\mathcal{E}_{\text{surf*}}\right|}{\left|\Delta\mathcal{E}_{\text{int*}}+\mathcal{E}_{\text{surf*}}\right|/2}.$ (12)
Because of the lack of a well-defined reference among the two quantities, their average $\left|\Delta\mathcal{E}_{\text{int*}}+\mathcal{E}_{\text{surf*}}\right|/2$ is used as normalisation factor [43].
Results
Quantification of the precision of FAHF
Time and space convergence
Following the guidelines of Sec. 3.3.2, time and space convergence are assessed. Table 1 and 2 respectively summarise the highlights of the two procedures, while Fig. 2 pictures the error behaviour.
Table 1: Input parameters of the time convergence procedure (blue font for those which vary), and resulting errors.
$\Delta x\;(\text{m})$ $N_{x}\;(-)$ $\Delta t\;(\text{s})$ $N_{t}\;(-)$ $\varepsilon_{*}(\Delta t)\;(-)$ $\varepsilon_{*}(\mathcal{E})\;(-)$
5.5E--04 60 4.0E--04 500 4.3E--05 3.07E--01
5.5E--04 60 1.0E--04 2000 1.0E--05 2.97E--01
5.5E--04 60 4.0E--05 5000 3.9E--06 2.95E--01
5.5E--04 60 2.0E--05 10000 1.7E--06 2.95E--01
5.5E--04 60 4.0E--06 50000 -- 2.94E--01
Table 2: Input parameters of the space convergence procedure (blue font for those which vary), and resulting errors. Reference FAHF simulation in azure (see text for details).
$\Delta x\;(\text{m})$ $N_{x}\;(-)$ $\Delta t\;(\text{s})$ $N_{t}\;(-)$ $\varepsilon_{*}(\Delta x)\;(-)$ $\varepsilon_{*}(\mathcal{E})\;(-)$
2.3E--03 15 7.2E--06 27777 2.1E--03 1.10E--00
1.1E--03 30 7.2E--06 27777 4.0E--04 5.63E--01
5.5E--04 60 7.2E--06 27777 7.1E--05 2.52E--01
2.7E--04 120 7.2E--06 27777 2.5E--05 1.18E--01
1.4E--04 240 7.2E--06 27777 7.1E--06 5.72E--02
6.8E--05 480 7.2E--06 27777 -- 2.81E--02
In particular, Fig. 2a shows how the ideal first-order convergence (black dashed line) is closely achieved by the error points $\varepsilon_{*}(\Delta t)$. Similarly, Fig. 2b confirms the second-order convergence of the error in space $\varepsilon_{*}(\Delta x)$, hence verifying the space convergence of the algorithms. The minor departures of the data-points from the ideal dashed lines are not of concern, and could be ascribed to a combination of errors arising from interpolation and boundary conditions (Sec. 3.1.4).
In conclusion, the combination $\Delta t=7.2\cdot 10^{-6}$ s and $\Delta x=\Delta y=1.4\cdot 10^{-4}$ m (azure in Table 2) is chosen as a reference for further analyses (Sec. 4.2). This choice guarantees a satisfactory precision in $T_{i,j}^{n}$, at an affordable computational cost (158 s per sub-chord).
Figure 2: (a) Time convergence and (b) space convergence of FAHF, showing the first-order and second-order decrease of the error evaluated as a function of $\Delta t$ and $\Delta x$, respectively.
Energy balance
According to Sec. 3.3.4, Fig. 3 depicts the behaviour of $\varepsilon_{*}(\mathcal{E})$ as a function of $\Delta x$, following from the data of which in Table 2 (last column).
The error $\varepsilon_{*}(\mathcal{E})$ is found to show a first-order convergence with $\Delta x=\Delta y$, and almost no dependence on $\Delta t$ (Table 1, last column). This is further investigated in Sec. 5.1. But because of the uncertainties customarily affecting experimental investigations of energy balance in tokamaks [51, 52], the error $\varepsilon_{*}(\mathcal{E})=5.72\%$ achieved according to Table 2 (azure) with the chosen $\Delta x=\Delta y=1.4\cdot 10^{-4}$ m and $\Delta t=7.2\cdot 10^{-6}$ s (Sec. 4.1.1) is suitable for the analyses.
It is further noted that the actual sign [^2] Recovered by removing the absolute value at the numerator of Eq. (12). of $\varepsilon_{*}(\mathcal{E})$ is such that $\Delta\mathcal{E}_{\text{int}}(t)>\mathcal{E}_{\text{surf}}(t)\;\forall t\in[40;140]$ ms if $\Delta x\geq 5.5\cdot 10^{-4}$ m. Instead, no well-defined sign of the error is present in the high-resolution cases $\Delta x\leq 2.7\cdot 10^{-4}$ m, as the two energies converge together.
Figure 3: Energy balance error $\varepsilon_{*}(\mathcal{E})$ as a function of $\Delta x$.
Surface heat flux: $\Delta x=\Delta y$
The surface heat flux resulting from the set of simulations of Sec. 4.1.1 (hence with $\Delta x=\Delta y$) is analysed according to Sec. 3.1.6, at time instant $t_{*}=128$ ms.
As found for the energy error in Sec. 4.1.2, $q_{i}^{n}$ does also not show a dependence on $\Delta t$, with the Eich fit parameters varying by 2% at most. Therefore, only the results from the simulations of space convergence in Table 2 are detailed in Table 3. Full $q_{i}^{n}$ profiles in space are then shown by the dots in Fig. 4a, with their corresponding fitted Eich functions (solid lines). The main fitting parameters are pictured in Fig. 4b.
On the one hand, a strong dependence on $\Delta x=\Delta y$ is evidenced, with the main fitting parameters in fact changing by up to a factor 2.5 as the resolution increases. On the other hand, a convergence of both the heat flux profiles and the corresponding fitting parameters is successfully verified as $\Delta x=\Delta y$ drops. Any imperfect convergence is ascribed to the simplified fitting procedure here adopted (Sec. 3.1.6).
The present section thus justifies how $\Delta t=7.2\cdot 10^{-6}$ s and $\Delta x=\Delta y=1.4\cdot 10^{-4}$ m (Sec. 4.1.1) are also suitable choices from the perspective of $q_{i}^{n}$, which is in close reach of the converged value (azure in Table 3).
Table 3: Eich fitting parameters [6] and Eich integral (Eq. (6)) as a function of $\Delta x=\Delta y$ in the simulated cases of Table 2. Reference FAHF simulation in azure (see text for details).
$\Delta x$ $N_{x}$ $\Delta y$ $N_{y}$ $q_{0}$ $\lambda_{q}$ $S$ $q_{\text{BG}}$ $x_{0}$ $f_{X}$ Integral
[…]
2.3E--03 15 2.3E--03 13 6.11E$+$06 5.35E--03 2.11E--03 2.53E$+$05 1.38E--03 3.28E$+$00 1.59E--01
1.1E--03 30 1.1E--03 26 9.16E$+$06 5.76E--03 1.70E--03 2.69E$+$05 2.72E--03 3.17E$+$00 2.23E--01
5.5E--04 60 5.5E--04 53 1.11E$+$07 6.68E--03 1.33E--03 2.45E$+$05 4.13E--03 2.87E$+$00 2.64E--01
2.7E--04 120 2.7E--04 107 1.19E$+$07 7.57E--03 9.48E--04 1.96E$+$05 4.98E--03 2.69E$+$00 2.83E--01
1.4E--04 240 1.4E--04 207 1.24E$+$07 1.04E--02 9.27E--04 3.50E$+$04 5.34E--03 2.10E$+$00 2.75E--01
6.8E--05 480 6.8E--05 426 1.26E$+$07 1.03E--02 7.81E--04 2.29E--08 5.52E--03 2.15E$+$00 2.78E--01
Figure 4: (a) Surface heat flux profiles computed by FAHF (dots), in the cases of Table 2 ($\Delta x=\Delta y$). Corresponding Eich fits in solid lines ($R^{2}>98.5\%$). (b) Main Eich fit parameters as a function of $\Delta x$.
Surface heat flux: $\Delta x\neq\Delta y$
The present paragraph revisits Sec. 4.1.3 in the light of $\Delta x\neq\Delta y$. For Eq. (5) depends on $\Delta y$ at the leading order, $\Delta x=5.5\cdot 10^{-4}$ m is here kept fixed, with $\Delta t=7.2\cdot 10^{-6}$ s (Table 2). Only $\Delta y$ is scanned according to Table 4 (first column), table which then reports the entire outcome of the scan at time $t_{*}=128$ ms. Heat flux profiles in space and main Eich fitting parameters are then depicted in Fig. 5.
The fitting parameters successfully converge towards fixed values, within a satisfactory 8% of their counterparts in Table 3 ($\Delta y=\Delta x$). This is however achieved at a diminished computational cost ($\propto N_{x}\cdot N_{y}\cdot N_{t}$): for instance, running FAHF with the $\Delta x=5.5\cdot 10^{-4}$ m and $\Delta y=6.3\cdot 10^{-5}$ m of Table 4 saves 45% of the computational time compared to the reference $\Delta x=\Delta y=1.4\cdot 10^{-4}$ m, for the same $\Delta t=7.2\cdot 10^{-6}$ s. If $\Delta t$ was instead chosen according to Eq. (3), almost one order of magnitude worth of save in computational time could be achieved.
Table 4: Eich fitting parameters [6] and Eich integral (Eq. (6)) as a function of $\Delta y\neq\Delta x$ with fixed $\Delta x=5.5\times 10^{-4}$ m and $\Delta t=7.2\times 10^{-6}$ s.
$\Delta x$ $N_{x}$ $\Delta y$ $N_{y}$ $q_{0}$ $\lambda_{q}$ $S$ $q_{\text{BG}}$ $x_{0}$ $f_{X}$ Integral
[…]
5.5E--04 60 2.0E--03 14 6.37E$+$06 5.84E--03 1.99E--03 2.34E$+$05 1.76E--03 3.07E$+$00 1.63E--01
5.5E--04 60 1.0E--03 29 9.17E$+$06 6.08E--03 1.67E--03 2.53E$+$05 3.01E--03 3.03E$+$00 2.21E--01
5.5E--04 60 5.0E--04 58 1.08E$+$07 6.45E--03 1.20E--03 2.34E$+$05 4.28E--03 2.99E$+$00 2.57E--01
5.5E--04 60 2.5E--04 116 1.15E$+$07 8.21E--03 9.77E--04 1.84E$+$05 5.05E--03 2.49E$+$00 2.73E--01
5.5E--04 60 1.3E--04 232 1.19E$+$07 1.01E--02 8.73E--04 1.78E$+$04 5.37E--03 2.17E$+$00 2.63E--01
5.5E--04 60 6.3E--05 464 1.21E$+$07 1.02E--02 7.65E--04 2.57E--12 5.53E--03 2.16E$+$00 2.67E--01
5.5E--04 60 3.1E--05 928 1.22E$+$07 1.03E--02 7.21E--04 4.47E--09 5.60E--03 2.16E$+$00 2.71E--01
Figure 5: (a) Surface heat flux profiles computed by FAHF (dots), in the cases of Table 4 ($\Delta y\neq\Delta x$). Corresponding Eich fits in solid lines ($R^{2}>98.5\%$). (b) Main Eich fit parameters as a function of $\Delta y$.
Estimate of the accuracy of FAHF: comparison against COMSOL Multiphysics®
Downstream proving the choice of $\Delta t=7.2\cdot 10^{-6}$ s and $\Delta x=\Delta y=1.4\cdot 10^{-4}$ m as acceptable in terms of convergence errors (Sec. 4.1.1), energy balance error (Sec. 4.1.2), and heat flux behaviour (Sec. 4.1.3), the necessary condition for a well-posed cross-code comparison among FAHF and COMSOL is ensured.
Assessment of the geometrical error in FAHF
According to Sec. 3.2, COMSOL is run in two (COMSOL2D) and three (COMSOL3D) dimensions, with temperature-independent material properties and with a single layer of molybdenum, hence matching FAHF models (Sec. 3.1.1). This allows insights in the geometrical error alone, as implied by Asm. #3.
The close agreement between FAHF, COMSOL2D and COMSOL3D is depicted by Fig. 6, which shows the comparison of: (a) the error $\varepsilon^{n}(q)$ in time according to Eq. (10); (b) the error $\varepsilon_{i}(q)$ in space from Eq. (11); (c) the maximum heat flux in time, $\text{max}_{i\in[1;N_{x}]}\{q_{i}^{n}\}$; (d) the heat flux profiles in space at time instant $t_{*}=128$ ms. The errors are computed with the FAHF result always taken as the reference $\tilde{q}$, while (c) and (d) offer the possibility of qualitative insights.
According to Fig. 6a, the satisfactory agreement between FAHF and COMSOL2D is evidenced for the entire time window $[40;140]\;\text{ms}$, with an error of $\varepsilon^{n}(q)=4.7\%\pm 1.3\%$. For FAHF vs. COMSOL3D, $\varepsilon^{n}(q)$ is clearly higher than its 2D counterpart, though still being an acceptable $8.3\%\pm 1.8\%$.
A significant inaccuracy of FAHF is observed when the plasma: is started ($t\sim 10$ ms); reverts back to being limited ($t\sim 160$ ms, hence suddenly depositing its power elsewhere than the camera FoV on the divertor); is terminated ($t\sim 200$ ms), i.e. when evolving on a short timescale [^3] The maximum heat flux fluctuations in $[100;150]$ ms are a combination between: saturation of $T_{\text{IR}}$ (Sec. 2.3); considering the only maximum of heat flux; plasma vertical oscillation. These do not fall under the umbrella of ’fast dynamics’. (Fig. 6c). Accurately capturing fast transients would indeed require proper numerical handling [14], but it is not of concern in the present study.
Fig. 6b suggests that similar conclusions hold for the error in space $\varepsilon_{i}(q)$, which satisfactorily attains values of $7.5\%\pm 1.4\%$ for FAHF vs. COMSOL2D all throughout the bulk of the tile, i.e. $x\in[2.5;30]$ mm. The enhanced error at the boundaries is imputable to FAHF there reverting to first-order BCs (Sec. 3.1.4). Slightly higher values of $10.0\%\pm 4.0\%$, though still acceptable, are found with respect to COMSOL3D.
Given the supportive evidence above, the appropriateness of adopting the multi-2D mode in FAHF in Asm. #3, rather than a full 3D treatment, is confirmed. This is also coherent with the findings of Ref. [46] concerning the multi-chord analysis.
Figure 6: (a) Error $\varepsilon^{n}(q)$ in time. (b) Error $\varepsilon_{i}(q)$ in space. (c) Maximum heat flux in time $\text{max}_{i\in[1;N_{x}]}\{q_{i}^{n}\}$. (d) Heat flux profiles in space at time instant $t_{*}=128$ ms over the reference sub-chord.
Assessment of the model errors in FAHF
Starting from the COMSOL2D simulation of Sec. 4.2.2 with (i) temperature-independent material properties and a single layer of 29-mm molybdenum (Mo()($\cdot$)( ⋅ )), the set of simulations is enlarged to assess the model simplifications (Asms. #1 and #2) of Sec. 3.1.1. COMSOL is indeed run in 2D: (ii) with temperature-dependent material properties [53, 54, 55] and a single layer of 29-mm molybdenum (Mo$(T)$); (iii) with temperature-independent material properties and a double layer of 4-mm molybdenum and 25-mm copper-cromium-zirconium (Mo+CuCrZr()($\cdot$)( ⋅ )); (iv) with temperature-dependent material properties [56] and a double layer of 4-mm molybdenum and 25-mm copper-cromium-zirconium (Mo+CuCrZr$(T)$).
FAHF, already compared against COMSOL2D in Sec. 4.2.1, is excluded from this comparison, to allow for the precise appreciation of the minute changes here involved, which would be obscured otherwise. The errors according to Eq. (10) and (11) are therefore computed with the COMSOL result Mo($\cdot$) taken as the reference $\tilde{q}$.
In particular, $\varepsilon^{n}(q)<0.5\%\;\forall t\in[15;150]$ ms in Fig. 7a, and $\varepsilon_{i}(q)<1\%\;\forall x$ in Fig. 7b. This confirms that Asms. #1 and #2 are appropriate in the present case.
Importantly, $\varepsilon^{n}(q)$ is practically zero if $t\lesssim 50$ ms for the case Mo($\cdot$) vs. Mo+CuCrZr($\cdot$). This confirms that any thermal information indeed takes tens of ms to reach the Mo-CuCrZr interface, located 4 mm in the depth of the divertor tiles, hence consistently with Sec. 3.1.1.
Also noticed in Sec. 4.2.1, the peculiar time windows of $t\lesssim 10$ ms and $t\gtrsim 160$ ms (of secondary importance in these analyses) violate the conclusions above, and are further commented in Sec. 5.4.1.
Figure 7: (a) Error $\varepsilon^{n}(q)$ in time and (b) error $\varepsilon_{i}(q)$ in space, computed among the two-dimensional COMSOL cases with respect to the COMSOL simulation implementing FAHF-like models.
Discussion
Convergence properties of the energy balance error
The goal of this paragraph is to explain the reasons behind the behaviour of $\varepsilon_{*}(\mathcal{E})$ of Sec. 4.1.2, with $\Delta x=\Delta y$. With the inequality $o(\Delta t)\ll o(\Delta x^{2})$ ensured in the space convergence (Sec. 3.3.2) and the convergence properties of Eq. (7) and Eq. (8), the error $\varepsilon_{*}(\mathcal{E})$ from Eq. (12) must abide to a first order convergence in space. This holds by virtue of the ordering $o(\Delta t)\ll o(\Delta x^{2})\sim o(\Delta x^{2})+o(\Delta y^{2})\ll o(\Delta y)\ll 1$.
On the contrary, an explicit dependence on $\Delta t$ is almost totally absent in Table 1 (last column). Granted $o(\Delta x^{2})\ll o(\Delta t)$ in the time convergence (Sec. 3.3.2), this suggests $o(\Delta x^{2})+o(\Delta y^{2})\sim o(\Delta x^{2})\ll o(\Delta t)\lesssim o(\Delta y)\ll 1$ to be in place, where $o(\Delta t)\lesssim o(\Delta y)$ implies that the trend in time of $\varepsilon_{*}(\mathcal{E})$ is hampered by $o(\Delta y)$.
In other words, the first-order error $o(\Delta y)$ affecting the heat flux $q_{i}^{n}$ via Eq. (5) is always the dominant contribution, with regards to energy conservation. Although the resulting error is still acceptable (Sec. 4.1.2), a point of improvement for FAHF would be using higher-order discretisations in the fashion of Ref. [49]’s instead of Eq. (5), were higher precision required.
Implications on choosing $\Delta x$ and $\Delta y$
As also suggested by Ref. [14], Sec. 4.1.4 clarifies how a precise heat flux can be calculated by the algorithm with a proper choice of $\Delta y<\Delta x$, without the need of an extremely small $\Delta x$. This hence allows for saving computational cost.
However, the assessment of the solution $q_{i}^{n}$ and its Eich fit in terms of $\Delta x$ (potentially with $\Delta x<\Delta x_{\text{IR}}$), as performed here, is also recommended. For small enough $\Delta y$, $\Delta x$ becomes the bottleneck in the error $o(\Delta x^{2})+o(\Delta y)$ of Eq. (5), even more so if higher-order discretisations with $\Delta y$ are employed.
Implications on choosing $\Delta x_{\text{IR}}$
The fact that the resolution along $y$ (independent on the diagnostic apparatus) mostly determines the precision of $q_{i}^{n}$ (within the limits of Sec. 5.2) does loosen the constraints on the required IR camera resolution $\Delta x_{\text{IR}}$ (Sec. 2.2).
The suitable $\Delta x_{\text{IR}}$ is such that to allow capturing features of $T_{\text{IR}}(x)$, a smoother and broader profile than its corresponding $q_{i}^{n}$. Instead, the suitable $\Delta x$ follows from Sec. 5.2. The two resulting figures can therefore differ, with $\Delta x_{\text{IR}}\geq\Delta x$, and potentially $\Delta x_{\text{IR}}\gg\Delta x$.
Nevertheless, the recurring diagnosis of power fall-off lengths significantly in defect of the Eich scaling predictions [57] does rise the concern on the resolution $\Delta x_{\text{IR}}$ required by IR systems in future devices.
Notably, the deviation of the FAHF result from the ideal Eich fit in $x\in[10;20]$ mm $\forall\Delta y\leq 1.3\cdot 10^{-4}$ m in Fig. 5a is a genuine feature, because $\Delta x=5.5\cdot 10^{-4}\;\text{m}$ adopted in Sec. 4.1.4 is above the IR camera resolution $\Delta x_{\text{IR}}=2.1\cdot 10^{-4}$ m (Sec. 2.2). The effective spatial resolution of the ST40 IR system does allow for the appreciation of fine-scale structures in the heat flux footprint, not dissimilar for those predicted by the XGC code[48, 58].
Geometrical error
Although Sec. 4.2.1 already concluded on the appropriateness of Asm. #3 in FAHF, further feature worth investigating appear from the comparison.
Features in time
The error FAHF vs. COMSOL3D of $9.6\%\pm 1.4\%$ in $[40;90]$ ms is greater than the $6.9\%\pm 1.1\%$ in $[90;140]$ ms by a statistically significant margin. The same is not true for FAHF vs. COMSOL2D instead, for which $4.6\%\pm 1.0\%$ in $[40;90]$ ms and $4.8\%\pm 1.6\%$ in $[90;140]$ ms. This 3D-related feature likely follows from the appearance of the strike point in the camera FoV during $[90;140]$ ms, which leads to $|\partial_{x}T_{\text{IR}}|\gg|\partial_{\phi}T_{\text{IR}}|$ (Sec. 3.1.2), i.e. a stronger tendency for the heat flux to flow poloidally (favouring a 2D treatment) rather than toroidally (inviting 3D flows).
Additionally, COMSOL3D maximum is always in defect of its COMSOL2D counterpart all throughout $[40;140]\;\text{ms}$. COMSOL3D hence predicts that the same PFC temperature (i.e. $T_{\text{IR}}$) tends to be reached by means of a lower surface heat flux. This effect is again tied to the three-dimensionality of the problem, and presumably due to the additional heat flux that flows along $\phi$ (i.e. across the chords, Sec. 3.1.2) from $k<10$ towards the reference chord $k=10$ in COMSOL3D (and in reality), by virtue of $\partial_{\phi}T_{\text{IR}}<0$ (Fig. 1b). This additional contribution along $\phi$, accounted neither by FAHF nor by COMSOL2D, helps rising the temperature along chord 10 in the 3D case, thus requiring a lower $q_{i}^{n}$ along $y$.
Moreover, because of the FAHF incapability of resolving rapid dynamics, FAHF errors diverge towards 100% during the fast transients at the instants $t\sim 10$ ms and $t\sim 160$ ms, over a time of $\sim 5$ ms. This is a thermal information which, in the legitimate approximation of semi-infinite solid [12] (Sec. 3.1.1), affects a region of thickness $2\sqrt{\alpha\cdot 5\;\text{ms}}\sim 1$ mm along $y$. According to the resistance-capacitor (RC) circuital interpretation[12], this information then exponentially dissipates with a time constant in the order of $(1\;\text{mm})^{2}/\alpha\sim 19$ ms. And the FAHF error is indeed seen to lose memory of the information over a timescale of this magnitude, following the $t\sim 10$ ms and $t\sim 160$ ms events. The exponential fit of the neat FAHF vs. COMSOL3D error over the time window $[10;35]$ ms returns a time constant of 12 ms with a satisfactory $R^{2}=98\%$. Notably, this implies that the error during the initial phase ($t\sim 10$ ms) does not detrimentally impact the longer-term dynamics predicted by FAHF.
Conversely, the error at $t\gtrsim 160$ ms, despite firstly dropping, remains significant. This is presumably implied by the nature of this phase, an inertial cooling after the plasma magnetic topology reverts back to limited (with a sudden drop of divertor heating). Such a simple dynamics of inertial cooling, especially in conjunction with adiabatic boundaries (Sec. 3.1.4), is highly sensitive to the conditions at its beginning ($t\sim 160$ ms), which are very different between FAHF and COMSOL ($\varepsilon^{n}(q)\sim 100$%).
Features in space
The FAHF vs. COMSOL3D error in space $\varepsilon_{i}(q)$ increases (i) towards the toroidal-running gaps, i.e. $x=0$ and $x=32$ mm. This is expected because of the first-order boundary conditions of FAHF (Sec. 3.1.4), and because of the higher degree of three-dimensionality necessarily involved around these regions: the adiabatic boundary invites the heat to flow elsewhere than along $x$, e.g. along $\phi$, not modelled in 2D. However, $\varepsilon_{i}(q)$ is also enhanced (ii) around the centre of the tile, i.e. $x\sim 15$ mm. This can be ascribed to the dual of (i): the centre of the tile allows for a 3D heat flux behaviour by offering the loosest constraints, by virtue of being the point farthest away from any adiabatic boundary. A gradual transition towards 2D between these 3D extremes (i) and (ii) instead characterises the remaining regions, where the error $\varepsilon_{i}(q)$ drops.
Negative heat fluxes
The detection of negative heat fluxes in IR thermographic inversions has been historically ascribed to (or a combination of) mechanical vibrations [16, 27], impurity surface layers [14, 15] and black-body radiation emission from the PFCs [15], at least.
Negative surface heat fluxes are here predicted in all the three cases: (i) in $(0;10]$ ms in Fig. 6c; (ii) at $t_{*}=128$ ms on the left of the peak in Fig. 6d, $x\in[0;5]$ mm; (iii) all throughout the the divertor cooling phase (Sec. 5.4.1) at $t\gtrsim 160$ ms, not shown.
With regards to (i), vibrations during the plasma start-up are the most reasonable explanation. The same does not instead hold for (ii), presumably affected by the presence of an impurity surface layer and non-negligible black-body emission. Last, the negative heat flux in (iii), comparable in magnitude to the maximum during such phase ($0.5\;\text{MW}\cdot\text{m}^{-2}$), is entirely imputable to the black-body emission during the inertial cooling. Both the above-mentioned points are worthwhile further investigating in the future, despite not impacting the conclusions of the present study.
Conclusions and outlook
Data of divertor surface temperature are recorded by ST40’s IR camera system with the highest effective spatial resolution available worldwide ($0.21\pm 0.03$ mm/pixel). Leveraging such highly-resolved data, FAHF, the new Tokamak Energy’s tool for inverse thermographic analyses, is developed as a finite-difference explicit solver of the unsteady heat conduction equation. Under the simplifying assumptions of temperature-independent material properties, single molybdenum layer instead of multi-material multi-layered PFCs, and two-dimensionality, FAHF calculates the surface heat flux impinging on the divertor.
Answering the question of whether FAHF qualifies as a suitable tool for inverse thermographic analyses is the focus of this work. And the answer is a positive one.
The formal verification of FAHF by means of time and space convergence is successful, in fact showing the expected first- and second-order convergence with $\Delta t$ and $\Delta x=\Delta y$ (for the sake of simplicity), respectively. The most appropriate combination of $\Delta t$ and $\Delta x$ is chosen accordingly, such as to guarantee a satisfactory compromise between precision of the temperature output and computational cost.
The verification is here further corroborated by an assessment of the energy balance. The error affecting the energy balance is evidenced to attain satisfactorily low values and: to feature a well-defined trend with $\Delta x$; to lack a dependence on $\Delta t$. Both such characteristics are successfully motivated in terms of ordering of the different contributions to the discretisation error, with $o(\Delta y)$ always being the dominant contribution. The energy-preserving properties of FAHF, whose assessment is recommended in the other codes, are hence deemed acceptable from the view point of the internal consistency of the code, and for applications to power balance studies in tokamaks.
The behaviour of the surface heat flux (along the divertor depth $y$), leading metric in any inverse thermographic analyses, is also thoroughly scrutinised. The Eich fit [6] is used to extract physical parameters of interest. The two separate cases of $\Delta x=\Delta y$ and $\Delta x\neq\Delta y$ are investigated. On the one hand, in both instances a strong sensitivity of the fitting parameters as a function of the spatial resolution is found, with variations up to a factor 2.5. On the other hand, the two instances ultimately show a convergence of the fitting parameters towards common values, as the spatial resolution is enhanced. Converged values of the $\Delta x\neq\Delta y$ case are achieved at a lower computational cost, by virtue of the leading dependence of the surface heat flux being $\Delta y$ (divertor depth), rather than $\Delta x$ (along the divertor). Although this dependence is commonly leveraged in inverse thermographic analyses [14, 49], the error implied by $\Delta x$ may become noticeable in the limit $\Delta y\ll\Delta x$, and detrimentally affect the surface heat flux. Routinely analysing the surface heat flux as a function of the space step $\Delta x$ along the divertor is thus recommended.
With the above guaranteeing a satisfactory precision in FAHF, the comparison with the renowned COMSOL Multiphysics® is well-posed, and evinces a satisfactory accuracy of FAHF in any instance analysed.
The average agreement within $8.3\%$ in time and $10\%$ in space among FAHF and COMSOL (3D) demonstrates the legitimacy of the 2D treatment adopted in FAHF. Furthermore, the simplistic implementation in FAHF of temperature-independent material properties and single molybdenum layer is proven to be appropriate in ST40 shots: the corrections implied by including these features in COMSOL (2D) are shown to be of minor entity ($<1\%$).
Regions of disagreement in time among FAHF and COMSOL are circumscribed to the beginning/end of the plasma shot, and at the termination of the divertor heating phase as the plasma abruptly reverts to being limited. These time windows of fast dynamics could be properly resolved by improving the temporal discretisation properties of FAHF [14], but are outside the scope of the present work, and outside the time window of interest for thermographic studies.
Disagreement in space among FAHF and COMSOL ($\sim 20\%$) is instead relegated to the regions in proximity of the toroidal-running gaps between divertor tiles. This is true in the 2D case, presumably because of the low-precision Neumann boundary conditions implemented in FAHF, and in the 3D case, also because of the adiabatic nature of these boundaries which invites a toroidal flow of the heat flux. Either instance does not pose a threat to the heat flux calculated by FAHF: the above-mentioned areas are also affected by further physical effects (e.g. tile side-heating due to finite ion Larmor radius [46]) which are not included in any thermographic inversion, and must therefore be under-emphasised anyway.
In conclusion, Tokamak Energy’s FAHF does qualify as a trustworthy tool, internally consistent and capable of providing a satisfactory precision and accuracy at an affordable computational cost, modulo a proper choice of $\Delta t$, $\Delta x$ and $\Delta y$ which this work identifies: $\Delta t=7.2\cdot 10^{-6}$ s and $\Delta x=\Delta y=1.4\cdot 10^{-4}$ m.
In perspective, further studies are planned in ST40 with the two-fold goal of improving the IR thermographic inversion experimentally and numerically. The current IR camera acquisition range has already been extended from $[20;118]$ to $[20;200]$, and a detailed cross-comparison with Langmuir probe measurements is underway. The installation of a second IR camera is being considered, for the exploitation of a dual-band adaptor [59] would allow ameliorating any dependence on the uncertain surface emissivity. The ongoing commissioning of the dedicated heat-flux testing facility in Tokamak Energy is ongoing, and will pave the way for a validation of FAHF in a controlled environment. Improvements to the numerics of FAHF to handle fast dynamics and impurity surface layers [14] are also taken into account.
Acknowledgments
This work was made possible by the US-DOE CRADA NFE-19-07769. Acknowledgment to the ever-precious help of David and Gabriela Hoffman, Sam Suchal, Teresa Thornton and Rosanne Monteiro is duely given.
Author declarations
The authors have no conflicts to disclose. The authors’ contribution statement using CRediT reads as follows. Matteo Moscheni: conceptualization (equal); methodology (equal); software (lead); formal analysis (lead); data curation (equal); visualization (lead); writing - original draft (lead); writing - review and editing (equal). Erik Maartensson: conceptualization (equal); methodology (equal); software (equal); writing - review and editing (equal). Matthew Robinson: software (equal); formal analysis (equal); writing - review and editing (equal). Chris Marsden: conceptualization (equal); software (equal); writing - review and editing (equal). Adrian Rengle: resources (lead); data curation (equal); writing - review and editing (equal). Travis Kelly Gray: conceptualization (equal); resources (equal); supervision (equal); writing - review and editing (equal). Andrea Scarabosio: supervision (equal); writing - review and editing (equal). Elena Vekshina: supervision (equal); writing - review and editing (equal). Xin Zhang: supervision (equal); writing - review and editing (equal).
Data availability statement
The data that support the findings of the study are available from the corresponding author upon reasonable request.
Appendix A: Calcam
The temperature information $T_{\text{IR}}$ from the IR camera view (Fig. 1b) belongs to the 2D pixel space, i.e. $T_{\text{IR}}=T_{\text{IR}}(x_{\text{p}},\phi_{\text{p}})$, with $x_{\text{p}}\in[1;540]$ and $\phi_{\text{p}}\in[1;612]$ (Sec. 2.2). During the pre-processing phase, the Calcam software [47] is employed to map such information in the three-dimensional CAD space ($\sim$ real space). The planar sub-space $x\phi$ of the CAD surface of the divertor tiles (Sec. 3.1.2) is ultimately of interest in FAHF. Calcam therefore creates the univocal relationship:
Calcam: $T_{\text{IR}}(x_{\text{p}},\phi_{\text{p}})\mapsto$ CAD space $\mapsto T_{\text{IR}}(x,\phi)$. (13)
The user input is required when building Eq. (13) via the Calcam GUI, and changes as a function of the camera location and alignment.
Appendix B: Python3 implementation
The entirety of the FAHF algorithms is written in Python3 [60], also because of its versatility when interfacing with third-party softwares (e.g. Calcam [47], Appendix A). Acknowledging the usual lack of implementation-specific information in the literature dealing with IR thermographical inversions, the actual implementation of Eq. (4) in Python3 is here concisely reported.
The meaning of the symbols is obvious, in fact coherent with the mathematical notation adopted throughout the paper. The only exception is for the indeces which are diminshed by 1, because of the Python3 convention of indexing from 0. {python}
import numpy as np from scipy.signal import convolve2d
# Fourier numbers for x and y Fo_x = alpha * dt / dx**2 Fo_y = alpha * dt / dy**2
# 5-point computational stencil stencil = np.array([ [ 0, Fo_y, 0 ], [ Fo_x, -2*Fo_x -2*Fo_y, Fo_x ], [ 0, Fo_y, 0 ] ])
# initialise temperature array: T = T(x,y,t) T = np.zeros((Nx, Ny, Nt))
# initial condition @n=0: T_IR = T_IR(x,t) T[:,:,0] = T_IR[:,0].mean()
for n in range(1,Nt+1):
# Dirichlet BC @j=0 T[:,0,n] = T_IR[:,n]
# explicitly compute temperature rise # mode = ’valid’ to avoid boundary nodes T_rise = convolve2d(in1 = T[:,:,n-1], in2 = stencil, mode = ’valid’)
# explicitly compute temperature @n # interior nodes only [1:-1,1:-1] T[1:-1,1:-1,n] = T[1:-1,1:-1,n-1] + T_rise
The code above exemplifies the implementation for interior and Dirichlet nodes. Neumann boundary nodes merely exploits a different shape of the stencil according to Ref. [12], and mode = ’same’ in convolve2d. Because of the linearity of the convolution convolve2d, the contribution of the boundary nodes can be separately added to T[:,:,n].
Notably, scipy’s function convolve2d is easily parallelised across multiple cores.